# Rethinking Approximate Gaussian Inference in Classification

### v1

@@ 6.1 Quality of Sample-Free Predictives
In this section, we investigate our first research question: whether there is a price to pay for sample-free predictives and second-order distributions. To this end, we take the two best-performing methods on the Expected Calibration Error (ECE) metric (Naeini et al., 2015), NormCDF and Sigmoid Laplace (c.f. Figure 6) and evaluate their per-predictive performance. As Table 1 shows, the analytic predictive is always on par with the (unbiased) MC estimates on ImageNet while being cheaper. This empirical observation supports our theoretical claim that when the constraint is satisfied, the analytic predictives are exact. Refer to Appendix L for an evaluation of the moment-matched Dirichlet distributions.
@@ 6.2 Effects of Changing the Learning Objective
The previous section shows that for an already trained model that satisfies the constraint (Equation 18), our analytic predictives always perform on par with MC sampling while being more efficient. However, our proposed objectives (Sections 4.1 and 4.2) change the training dynamics of models. Therefore, in this section, we showcase the performance of methods equipped with our analytic (i.e., sample-free) predictives and learning objectives against softmax models. For the latter, we use the best-performing predictive and estimator (excluding ours). See Appendices F and G for an overview of predictives and estimators. We employ our methods with the analytic predictives and second-order Dirichlet distributions to demonstrate their competitive nature while being more efficient than MC sampling.
Figure 4: ImageNet log probability proper scoring results. Analytic predictives with implicit ( ) constraints are either on par with or outperform Softmax ( ) on all methods.
Figure 5: CIFAR-10 log probability proper scoring results. Analytic predictives with implicit ( ) constraints consistently outperform Softmax ( ) on all methods.
@@
We first evaluate how calibrated the models are using the log probability scoring rule (Gneiting & Raftery, 2007) and the Expected Calibration Error (ECE) metric (Naeini et al., 2015). Figure 5 shows that on CIFAR-10, the score of our analytic predictives with implicit constraints (Sigmoid, NormCDF) are consistently better than the corresponding Softmax results for all methods. On the large-scale ImageNet dataset, Sigmoid and NormCDF either outperform or are on par with Softmax.
Figure 6: ImageNet ECE results. Analytic predictives with implicit constraints ( ) outperform Softmax ( ) across all methods. Explicit constraints ( ) work best on HET but do not benefit the other two methods. Note the restricted $y$-limits for readability.
The ECE requires the models’ confidence to match their accuracy. Figure 6 shows that on ImageNet, our analytic predictives with implicit constraints (Sigmoid, NormCDF) have a clear advantage over Softmax across all methods. Explicit constraints benefit SNGP and HET but not the post-hoc Laplace method.
@@
Figure 7: ImageNet log probability proper scoring results for the binary correctness prediction task. Analytic predictives with implicit ( ) constraints consistently outperform Softmax ( ).
@@
Next, we turn to the correctness prediction task: whether the models can predict the correctness of their own predictions. We consider correctness estimators $\tilde{c}(x)\in[0,1]$ for inputs $x\in\mathcal{X}$ derived from the predictives. Framed as a binary prediction task, the goal of these estimators is to predict the probability of the predicted class’ correctness. We then measure the binary log probability score of $\tilde{c}(x)$. Figure 7 shows that our analytic predictives with implicit constraints outperform all Softmax predictives across all methods.
@@
Figure 8: Analytic predictives do not sacrifice accuracy. ImageNet validation accuracies. Analytic predictives with implicit ( ) or explicit ( ) constraints either outperform or are on par with Softmax ( ) across all methods.
Analytic predictives do not sacrifice accuracy.
@@
Figure 8 evidences this claim on ImageNet: our analytic predictives either outperform or are on par with Softmax predictives. The most accurate methods are NormCDF and Sigmoid SNGP. These results are in line with the findings of Wightman et al. (2021), who also recommend training with a per-class binary cross-entropy loss using the sigmoid activation function.
Figure 9: ImageNet OOD detection AUROC results for severity level one. Across all methods, the best-performing predictive is analytic with an implicit constraint ( ).
@@
Finally, we consider another binary prediction task, where a general uncertainty estimator $u(x)\in\mathbb{R}$ (derived from predictives or second-order Dirichlet distributions) is tasked to separate ID and OOD samples from a balanced mixture. As the uncertainty estimator can take on any real value, we measure the Area Under the Receiver Operating Characteristic curve (AUROC), which quantifies the separability of ID and OOD samples w.r.t. the uncertainty estimator. As OOD inputs, we consider corrupted ImageNet-C (Hendrycks & Dietterich, 2019) samples. Figure 9 shows that our analytic predictives with implicit constraints (Sigmoid, NormCDF) outperform Softmax across all methods, even though separating ID and OOD samples does not require a fine-grained representation of uncertainty, unlike the ECE or proper scoring rules. Importantly, these analytic predictives and second-order Dirichlet distributions are considerably cheaper to calculate than Softmax MC predictions (see Section 5). See Appendix M for the other severity levels.
@@ 7 Conclusion and Limitations
We developed a framework that allows to obtain predictives and other quantities of interest from logit space Gaussian distributions analytically. Our experimental results suggest the ubiquitous softmax activation should be replaced by normCDF or sigmoid.
We proposed to approximate the Gaussian pushforwards by Dirichlet distributions which cannot encode correlations between classes. Leveraging a more expressive yet tractable family of distributions on the simplex that can achieve this presents an interesting area of future research.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Gaussian Integral Derivations
In this appendix, we derive the closed-form formula for the mean of Gaussian pushforwards through exp (Equation 8) and normCDF (Equation 14), as well as the approximations for pushforwards through sigmoid (Equation 16) and softmax.
@@ A.2 Gaussian NormCDF Integral
Here, we derive the classical normCDF Gaussian integration formula (Owen, 1980, Eq. 10,010.8).
@@ A.3 Probit Approximation for Gaussian Sigmoid Integral
Hence matching $\rho$ and $\Phi$ to first order we get the approximation $\rho(y)\approx\Phi\left(\sqrt{\frac{\pi}{8}}y\right)$. So using Equation 33 we derive the probit approximation (Spiegelhalter & Lauritzen, 1990; MacKay, 1992)
@@ A.4 Mean Field Approximation for Gaussian Softmax Integral
For $\bm{\mu}\in\mathbb{R}^{C}$, $\bm{\sigma}^{2}\in\mathbb{R}^{C}_{>0}$ and $\bm{\Sigma}=\operatorname{\mathbf{diag}}(\bm{\sigma}^{2})$, the mean field approximation to the Gaussian softmax integral (Lu et al., 2021) is obtained as follows
@@
(1) is the mean field approximation, and (3) uses the probit approximation Equation 35. (Lu et al., 2021) provides two other variants of this approximation with other choices of approximation (2).
@@ Appendix B Comparison with the Multinomial Probit Model
In this appendix, we show that a model whose output activation is a composition of an element-wise normCDF activation $\bm{\Phi}$ and a normalisation $\bm{n}$ is distinct from the classical multinomial probit model (Daganzo, 1979).
@@
In the case $C=2$, the multinomial probit model Equation 39 outputs closed form probabilities. This allows us to construct an explicit counterexample to the equivalence of the two models Equation 39 and Equation 40:
@@
Appendix C Analysis of the Predictive Approximations’ Quality
In this appendix, we provide theoretical analyses of the quality of various predictive approximations. This complements the empirical analyses taken, for instance, in the synthetic experiment (Figure 1) or in (Daunizeau, 2017).
Due to its information-theoretic interpretation, a natural metric to equip the probability simplex $\Delta^{C-1}$ with is the KL divergence
@@ Appendix C Analysis of the Predictive Approximations’ Quality
C.1 Analysis of Monte Carlo Approximations
@@
We now give an informal theoretical argument for the linear growth of the KL divergence between $\bm{p}$ and $\bm{\hat{P}}^{\lceil S/C\rceil}$ in terms of $C$, under the distributional conditions of the synthetic experiment (Figure 1).
Taylor expanding Equation 42 about $\bm{p}$ to second order we obtain
@@
$=\sum_{c=1}^{C}p_{c}\sum_{k=1}^{\infty}\frac{(p_{c}-q_{c})^{k}}{kp_{c}^{k}}$
$=\underbrace{\sum_{c=1}^{C}p_{c}}_{=1}-\underbrace{\sum_{c=1}^{C}q_{c}}_{=1}+\sum_{c=1}^{C}\sum_{k=2}^{\infty}\frac{(p_{c}-q_{c})^{k}}{kp_{c}^{k-1}}$
$\stackrel{{\scriptstyle(1)}}{{\approx}}\sum_{c=1}^{C}\frac{(p_{c}-q_{c})^{2}}{2p_{c}}$
@@
In the synthetic experiment (Figure 1), the logit class-wise means $\mu_{c}$ and variances $\sigma_{c}^{2}$ are sampled in an i.i.d. way. Let $\bm{Y}\sim N(\bm{\mu},\bm{\Sigma})$ with $\bm{\Sigma}:=\operatorname{\mathbf{diag}}(\bm{\sigma}^{2})$, $\bm{Q}=\bm{\varphi}(\bm{Y})$ the unnormalised ‘probabilities’ and $\bm{P}:=\bm{Q}/\sum_{c=1}^{C}Q_{c}$ the probabilities. We have
@@
Plugging this into Equation 44 we get
@@
which grows linearly with the number of classes, as observed in Figure 1.
C.2 Analysis of the Probit Approximation for Sigmoid Models
Recall the probit approximation to the Gaussian sigmoid integral (Section A.3)
$q:=\int_{\mathbb{R}}\rho(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)\approx\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)=:\hat{q}.$ (49)
In the multiclass setting, for $\bm{\mu}\in\mathbb{R}^{C}$, $\bm{\sigma}^{2}\in\mathbb{R}^{C}_{>0}$ and $\bm{\Sigma}=\operatorname{\mathbf{diag}}(\bm{\sigma}^{2})$, define
$\bm{q}:=\mathbb{E}$ [${}_{\bm{Q}\sim\bm{\rho}_{*}\mathcal{N}(\bm{\mu},\bm{\Sigma})}[\bm{Q}]$ (50)
$\bm{\hat{q}}:=\bm{\rho}$ $\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right)$
$\bm{p}:=\mathbb{E}_{\bm{P}\sim\bm{n}_{*}\bm{\rho}_{*}\mathcal{N}(\bm{\mu},\bm{\Sigma})}$ $[\bm{P}]=\mathbb{E}_{\bm{Q}\sim\bm{\rho}_{*}\mathcal{N}(\bm{\mu},\bm{\Sigma})}\left[\frac{\bm{Q}}{\sum_{c=1}^{C}Q_{c}}\right]$
$\bm{\hat{p}}:=\frac{\bm{\hat{q}}}{\sum_{c=1}^{C}\hat{q}_{c}}$ $=\frac{\bm{\rho}\left(\frac{\bm{\mu}}{\sqrt{1+\frac{\pi}{8}\bm{\sigma}^{2}}}\right)}{\sum_{c=1}^{C}\rho\left(\frac{\mu_{c}}{\sqrt{1+\frac{\pi}{8}\sigma^{2}_{c}}}\right)}$
We assume that, e.g. through training, the constraint Equation 18 is perfectly satisfied, such that $\bm{p}=\bm{q}/\sum_{c=1}^{C}q_{c}$. Then we obtain
@@ C.2 Analysis of the Probit Approximation for Sigmoid Models
Let $\mathcal{K}\subset\mathbb{R}\times\mathbb{R}_{>0}$ compact. Using the compactness of $\mathcal{K}$ define $\delta(\mathcal{K}):=\max_{(\mu,\sigma^{2})\in\mathcal{K}}|q-\hat{q}|$ and $u(\mathcal{K}):=\min_{(\mu,\sigma^{2})\in\mathcal{K}}q>0$, where $q$ and $\hat{q}$ are as in Equation 49. Then if $(\mu_{c},\sigma_{c}^{2})\in\mathcal{K}$ for all $c$ we have
$\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})\leq\log\left(\frac{u+\delta}{u-\delta}\right).$ (51)
Remark C.2.
The bound Equation 51 is of practical value as

### v2

@@ 6.1 Quality of Sample-Free Predictives
In this section, we investigate whether there is a price to pay for sample-free predictives and second-order distributions. To this end, we take the two best-performing methods on the Expected Calibration Error (ECE) metric [29]—NormCDF and Sigmoid Laplace (see Fig. 4(a))—and evaluate their per-predictive performance. As Table 1 shows, the closed-form predictive is always on par with the (unbiased) MC estimates on ImageNet while being cheaper. Section K.4 reports the KL divergence of different approximations to the ‘true’ predictive (estimated using 10,000 MC samples) on ImageNet. Table 2 shows that the moment-matched second-order Dirichlet distributions consistently outperform both MC and sample-free predictives on the OOD detection task, showcasing the practical utility of second-order distributional knowledge.
(a) Closed-form predictives ( ) are either on par with or outperform Softmax ( ) across all methods.
(b) Closed-form predictives ( ) consistently outperform Softmax ( ) across all methods.
Figure 3: ImageNet (left) and CIFAR-10 (right) test negative log likelihood results.
@@ 6.2 Effects of Changing the Learning Objective
The previous section shows that for models trained with Sigmoid or NormCDF, our closed-form predictives always perform on par with MC sampling while being more efficient. However, our proposed objective (see Eq. 29) changes the training dynamics and, subsequently, the performance of models. In this section, we showcase the performance of methods equipped with our closed-form predictives and learning objectives against Softmax models. For the latter, we use the best-performing predictive and estimator (see Appendices E and F for an overview of predictive approximations and estimators). We employ our methods with the closed-form predictives and second-order Dirichlet distributions to demonstrate their competitive nature while being more efficient than MC sampling.
@@
We first evaluate how calibrated the models are using the test NLL [12] and Expected Calibration Error (ECE) [29] metrics. Fig. 3(b) shows that on CIFAR-10, the score of our closed-form predictives (Sigmoid, NormCDF) is consistently better than the corresponding Softmax results for all methods. On the large-scale ImageNet dataset, Sigmoid and NormCDF predictives either outperform or are on par with Softmax (see Fig. 3(a)).
(a) Closed-form predictives ( ) are more calibrated than Softmax ( ) across all methods. Note the restricted $y$-limits for readability.
(b) Our closed-form predictives ( ) consistently outperform Softmax ( ) on predicting their own correctness.
@@
Figure 4: ImageNet ECE (left) and log probability proper scoring results for the binary correctness prediction task (right).
The ECE metric requires the models’ confidence (i.e., maximum class probability) to match their accuracy. Fig. 4(a) shows that, on ImageNet, our closed-form predictives have a clear advantage over Softmax across all methods.
@@
Next, we turn to the correctness prediction task: whether the models can predict the correctness of their own predictions. We consider correctness estimators $t(x):=\max_{1\leq c\leq C}\mathbb{E}_{\mathbf{P}\sim\mathrm{h}(x)}\left[P_{c}\right]\in[0,1]$ for inputs $x\in\mathcal{X}$ derived from the predictive mean $\mathbb{E}_{\mathbf{P}\sim\mathrm{h}(x)}\left[\mathbf{P}\right]$. Framed as a binary prediction task, the goal of these estimators is to predict the probability of the predicted class’ correctness, defined as $v(x)=[\hat{y}(x)=y]\in\{0,1\}$ with $\hat{y}(x):=\arg\max_{1\leq c\leq C}\mathbb{E}_{\mathbf{P}\sim\mathrm{h}(x)}\left[P_{c}\right]\in\{1,\dots,C\}$. We then measure the binary log probability score of $\tilde{t}(x)$, given by $v(x)\log t(x)+(1-v(x))\log(1-t(x))$. Fig. 4(b) shows that our closed-form predictives outperform all Softmax predictives across all methods.
@@
(a) Closed-form predictives do not sacrifice accuracy. Closed-form predictives ( ) either outperform or are on par with Softmax ( ) across all methods.
(b) Closed-form predictives yield superior OOD detection performance. Across all methods, the best-performing predictive is closed-form ( ).
Figure 5: ImageNet validation accuracy (left) and OOD detection AUROC results for ImageNet-C transforms of severity level one (right).
Closed-form predictives do not sacrifice accuracy.
@@
Fig. 5(a) evidences this claim on ImageNet: our closed-form predictives either outperform or are on par with Softmax predictives. These results are in line with the findings of Wightman et al. [38], who also recommend training with a per-class binary cross-entropy loss using the sigmoid activation function. Section K.5 shows that the BCE loss already yields performance gains in the model backbones (without the logit-space Gaussians).
@@
Finally, we consider another binary prediction task, where a general uncertainty estimator $u(x)\in\mathbb{R}$ (derived from predictives or second-order Dirichlet distributions) is tasked to separate ID and OOD samples from a balanced mixture thereof. As the uncertainty estimator can take on any real value, we measure the Area Under the Receiver Operating Characteristic curve (AUROC), which quantifies the separability of ID and OOD samples w.r.t. the uncertainty estimator. As OOD inputs, we consider corrupted ImageNet-C [13] samples. Fig. 5(b) shows that our closed-form predictives outperform Softmax across all methods. Importantly, these closed-form predictives and second-order Dirichlet distributions are considerably cheaper to calculate than Softmax MC predictions (see Section 5). See Section K.6 for the other severity levels.
@@ 7 Conclusion and Limitations
We developed a framework that allows for obtaining predictives and other quantities of interest from logit space Gaussian distributions in closed form. Our experimental results suggest that the ubiquitous softmax activation should be replaced by normCDF or sigmoid for uncertainty quantification tasks.
A limitation is that our approximate predictives and approximate Dirichlet distributions do not encode correlations between classes. While our theoretical analysis suggests strong bounds on the approximations, leveraging more expressive yet tractable second-order distributions on the simplex that can model correlations presents an interesting area of future research.
@@ Appendix A Gaussian Integral Derivations
In this appendix, we derive the closed-form formula for the mean of Gaussian pushforwards through exp (Eq. 8) and normCDF (Eq. 13), as well as the approximations for pushforwards through sigmoid (Eq. 15) and softmax.
@@ A.2 Gaussian NormCDF Integral
Here, we derive the classical normCDF Gaussian integration formula [30, Eq. 10,010.8].
@@ A.3 Probit Approximation for Gaussian Sigmoid Integral
as $y\to 0$. Hence matching $\rho$ and $\Phi$ to first order we get the approximation $\rho(y)\approx\Phi\left(\sqrt{\frac{\pi}{8}}y\right)$. So using Eq. 31 we derive the probit approximation [34, 23]
@@ A.4 Mean Field Approximation for Gaussian Softmax Integral
For $\bm{\mu}\in\mathbb{R}^{C}$, $\bm{\sigma}^{2}\in\mathbb{R}^{C}_{>0}$ and $\bm{\Sigma}=\operatorname{\mathbf{diag}}(\bm{\sigma}^{2})$, the mean field approximation to the Gaussian softmax integral [21] is obtained as follows
@@
(1) is the mean field approximation, and (3) uses the probit approximation Eq. 33. [21] provides two other variants of this approximation with other choices of approximation (2).
@@ Appendix B Comparison with the Multinomial Probit Model
In this appendix, we show that a model whose output activation is a composition of an element-wise normCDF activation $\bm{\Phi}$ and a normalisation $\bm{n}$ is distinct from the classical multinomial probit model [6].
@@
In the case $C=2$, the multinomial probit model Eq. 37 outputs closed-form probabilities. This allows us to construct an explicit counterexample to the equivalence of the two models Eq. 37 and Eq. 38:
@@
Appendix C Theoretical Analyses
In this appendix, we provide (formal and informal) theoretical analyses of the quality of various predictive approximations. This complements the empirical analyses, for instance in the synthetic experiment (Fig. 1) or in [7].
Due to its information-theoretic interpretation, a natural divergence to equip the probability simplex $\Delta^{C-1}$ with is the Kullback-Leibler (KL) divergence
@@ Appendix C Theoretical Analyses
C.1 Informal Analysis of Monte Carlo Approximations
@@
We now give an informal theoretical argument for the linear growth of the KL divergence between $\bm{p}$ and $\bm{\hat{P}}^{\lceil S/C\rceil}$ in terms of $C$, under the distributional conditions of the synthetic experiment (Fig. 1).
Taylor expanding Eq. 40 about $\bm{p}$ to second order we obtain
@@
$\stackrel{{\scriptstyle(1)}}{{\approx}}\sum_{c=1}^{C}p_{c}\left(1-\frac{p_{c}}{q_{c}}+\frac{(p_{c}-q_{c})^{2}}{2p_{c}^{2}}\right)$
$=\underbrace{\sum_{c=1}^{C}p_{c}}_{=1}-\underbrace{\sum_{c=1}^{C}q_{c}}_{=1}+\sum_{c=1}^{C}\frac{(p_{c}-q_{c})^{2}}{2p_{c}}$
$=\sum_{c=1}^{C}\frac{(p_{c}-q_{c})^{2}}{2p_{c}}$
@@
In the synthetic experiment (Fig. 1), the logit class-wise means $\mu_{c}$ and variances $\sigma_{c}^{2}$ are sampled in an i.i.d. way. Let $\bm{Q}\sim\mathcal{N}(\bm{\mu},\operatorname{\mathbf{diag}}(\bm{\sigma}^{2}))$ the unnormalised ‘probabilities’ and $\bm{P}:=\bm{Q}/\sum_{c=1}^{C}Q_{c}$ the probabilities. We have
@@
Plugging this into Eq. 42 we get
@@
which grows linearly with the number of classes, as observed in Fig. 1.
C.2 Analysis of the Closed-Form Approximations (Theorem 2.1)
In this section we prove Theorem 2.1 and discuss whether its underlying assumptions are fulfilled in practice. We start with a lemma.
Lemma C.1.
Let $(X_{n})_{n\geq 1},(Y_{n})_{n\geq 1}$ be sequences of $(0,M)$-valued random variables for some $M>0$. Suppose $\sup_{n}\mathbb{E}[Y_{n}^{-8}]<\infty$ and that $\mathrm{Var}(Y_{n})\to 0$ as $n\to\infty$. Then
$\left|\mathbb{E}\left[\frac{X_{n}}{Y_{n}}\right]-\frac{\mathbb{E}[X_{n}]}{\mathbb{E}[Y_{n}]}\right|=O(\mathrm{Var}(Y_{n})^{1/2})$ (47)
as $n\to\infty$.
Proof.
Let $X,Y$ be $(0,M)$-valued random variables. We take a bivariate Taylor expansion of $\frac{X}{Y}$ for the random variables $X$ and $Y$ around $\mathbb{E}[X]$ and $\mathbb{E}[Y]$ respectively:
$\frac{X}{Y}=$ $\frac{\mathbb{E}[X]}{\mathbb{E}[Y]}+\frac{1}{\mathbb{E}[Y]}(X-\mathbb{E}[X])-\frac{\mathbb{E}[X]}{\mathbb{E}[Y]^{2}}(Y-\mathbb{E}[Y])$ (48)
$-\frac{1}{\xi(Y)^{2}}(X-\mathbb{E}[X])(Y-\mathbb{E}[Y])+\frac{\eta(X)}{\xi(Y)^{3}}(Y-\mathbb{E}[Y])^{2}$
where $\eta(X)\in[\mathbb{E}[X],X]\text{ or }[X,\mathbb{E}[X]]$ and $\xi(Y)\in[\mathbb{E}[Y],Y]\text{ or }[Y,\mathbb{E}[Y]]$, using the Lagrange form of the remainder. So by the Cauchy-Schwarz inequality,
$\left|\mathbb{E}\left[\frac{X}{Y}\right]-\frac{\mathbb{E}[X]}{\mathbb{E}[Y]}\right|$ $=\left|-\mathbb{E}\left[\frac{1}{\xi(Y)^{2}}(X-\mathbb{E}[X])(Y-\mathbb{E}[Y])\right]+\mathbb{E}\left[\frac{\eta(X)}{\xi(Y)^{3}}(Y-\mathbb{E}[Y])^{2}\right]\right|$ (49)
$\leq\mathbb{E}\left[\frac{1}{\xi(Y)^{8}}\right]^{1/4}\mathbb{E}[(X-\mathbb{E}[X])^{4}]^{1/4}\mathbb{E}[(Y-\mathbb{E}[Y])^{2}]^{1/2}$
$\quad+\mathbb{E}\left[\frac{\eta(X)^{2}}{\xi(Y)^{6}}\right]^{1/2}\mathbb{E}[(Y-\mathbb{E}[Y])^{4}]^{1/2}.$
Note that
$\eta(X_{n})$ $\leq M,$ (50)
$\sup_{n}\mathbb{E}\left[\frac{1}{\xi(Y_{n})^{k}}\right]$ $\leq\sup_{n}\mathbb{E}\left[\frac{1}{\min(Y_{n}^{k},\mathbb{E}[Y_{n}]^{k})}\right]\leq\sup_{n}\mathbb{E}\left[\frac{1}{Y_{n}^{k}}+\frac{1}{\mathbb{E}[Y_{n}]^{k}}\right]<\infty\text{ for }k\in\{6,8\},$
$\mathbb{E}[(X_{n}-\mathbb{E}[X_{n}])^{4}]$ $\leq M^{4},$
$\mathbb{E}[(Y_{n}-\mathbb{E}[Y_{n}])^{4}]$ $\leq M^{2}\mathbb{E}[(Y_{n}-\mathbb{E}[Y_{n}])^{2}].$
So from Eq. 49 we obtain
$\left|\mathbb{E}\left[\frac{X_{n}}{Y_{n}}\right]-\frac{\mathbb{E}[X_{n}]}{\mathbb{E}[Y_{n}]}\right|=O(\mathbb{E}[(Y_{n}-\mathbb{E}[Y_{n}])^{2}]^{1/2})=O(\mathrm{Var}(Y_{n})^{1/2})$ (51)
as $n\to\infty$. ∎
Now recall our notation
$q=q(\mu,\sigma^{2}):=\mathbb{E}_{Q\sim\rho_{*}\mathcal{N}(\mu,\sigma^{2})}[Q]\approx\rho\left(\frac{\mu}{\sqrt{1+\frac{\pi}{8}\sigma^{2}}}\right)=:\hat{q}(\mu,\sigma^{2})=\hat{q}$ (52)
where $\hat{q}$ is the probit approximation (Section A.3). Moreover we write $q_{c}:=q(\mu_{c},\sigma_{c}^{2})$, $\hat{q}_{c}:=\hat{q}(\mu_{c},\sigma_{c}^{2})$, $\bm{p}$ for the true predictive, $\bm{\hat{p}}$ for our approximate predictive, and in addition $\bm{\tilde{p}}$ for the approximate predictive using the exact one dimensional integrals, i.e.
$\bm{p}:=\mathbb{E}_{\bm{P}\sim\bm{a}_{*}\mathcal{N}(\bm{\mu},\bm{\Sigma})}[\bm{P}],\quad\bm{\hat{p}}:=\frac{\hat{q}_{c}}{\sum_{c^{\prime}=1}^{C}\hat{q}_{c^{\prime}}},\quad\bm{\tilde{p}}:=\frac{q_{c}}{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}}.$ (53)
We now restate Theorem 2.1 with an explicit expression for $M(\mathcal{K})$:
@@
Suppose the means and variances $(\mu_{c},\sigma_{c}^{2})$ lie in some compact set $\mathcal{K}\subset\mathbb{R}\times[0,\infty)$ for each class $c$. Using the compactness of $\mathcal{K}$, define
  • $\delta(\mathcal{K}):=\sup_{(\mu,\sigma^{2})\in\mathcal{K}}(\hat{q}-q)$,
  • $u(\mathcal{K}):=\inf_{(\mu,\sigma^{2})\in\mathcal{K}}q>0$,
  • $\Delta(\mathcal{K}):=\sup_{(\mu,\sigma^{2})\in\mathcal{K}}\frac{q-\hat{q}}{q}$.
If $\Delta(\mathcal{K})>1$ then
$\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})\leq\log\left(\frac{1+\delta/u}{1-\Delta}\right)+O\left(\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\right)$ (54)
as $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$
Proof.
We have
$\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})=\sum_{c=1}^{C}p_{c}(\log p_{c}-\log\hat{p}_{c})=\underbrace{\sum_{c=1}^{C}p_{c}(\log p_{c}-\log\tilde{p}_{c})}_{(1)}+\underbrace{\sum_{c=1}^{C}p_{c}(\log\tilde{p}_{c}-\log\hat{p}_{c})}_{(2)}.$ (55)
Assuming $\Delta<1$, we can bound (2):
$(2)$ $=\sum_{c=1}^{C}\frac{q_{c}}{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}}\left(\log\left(\frac{q_{c}}{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}}\right)-\log\left(\frac{\hat{q}_{c}}{\sum_{c^{\prime}=1}^{C}\hat{q}_{c^{\prime}}}\right)\right)$ (56)
$=\frac{1}{\sum_{c=1}^{C}q_{c}}\sum_{c=1}^{C}q_{c}\left(-\log\left(1-\frac{q_{c}-\hat{q}_{c}}{q_{c}}\right)+\log\left(1+\frac{\sum_{c^{\prime}=1}^{C}\hat{q}_{c^{\prime}}-\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}}{\sum_{c^{\prime}=1}^{C}q_{c^{\prime}}}\right)\right)$
$\leq\frac{1}{\sum_{c=1}^{C}q_{c}}\sum_{c=1}^{C}q_{c}\left(-\log\left(1-\Delta\right)+\log\left(1+\frac{C\delta}{Cu}\right)\right)$
$=\frac{1}{\sum_{c=1}^{C}q_{c}}\sum_{c=1}^{C}q_{c}\log\left(\frac{1+\delta/u}{1-\Delta}\right)$
$=\log\left(\frac{1+\delta/u}{1-\Delta}\right).$
Now for (1), first note that $\mathbb{E}\left[\left(\sum_{c=1}^{C}Q_{c}\right)^{-8}\right]<\infty$ where [^2] In the case $\varphi=\Phi$ and $Q_{c}\sim\Phi_{*}\mathcal{N}(\mu_{c},\sigma_{c}^{2})$, due to the fast decay of the tail of $\Phi$ we may have $\mathbb{E}\left[\left(\sum_{c=1}^{C}Q_{c}\right)^{-8}\right]=\infty$, making the proof strategy fail in that case. $Q_{c}\sim\rho_{*}\mathcal{N}(\mu_{c},\sigma_{c}^{2})$. By compactness of $\mathcal{K}$, we have in fact $\sup_{(\bm{\mu},\bm{\sigma})\in\mathcal{K}^{C}}\mathbb{E}\left[\left(\sum_{c=1}^{C}Q_{c}\right)^{-8}\right]<\infty$. Thus we can apply Lemma C.1 with $(X_{n})_{n\geq 1}$ a sequence of $Q_{c}\sim\rho_{*}\mathcal{N}(\mu_{c},\sigma_{c}^{2})$ and $(Y_{n})_{n\geq 1}$ a sequence of $\sum_{c=1}^{C}Q_{c}$, to get
$|p_{c}-\tilde{p}_{c}|=O\left(\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)^{1/2}\right)$ (57)
as $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$. Thus Taylor expanding each term of (1) around $p_{c}$ we get
$(1)$ $=\sum_{c=1}^{C}p_{c}\left(\frac{p_{c}-\tilde{p}_{c}}{p_{c}}+\frac{(p_{c}-\tilde{p}_{c})^{2}}{2\omega(\tilde{p}_{c})^{2}}\right)$ (58)
$=\underbrace{\sum_{c=1}^{C}p_{c}}_{=1}-\underbrace{\sum_{c=1}^{C}\tilde{p}_{c}}_{=1}+\frac{(p_{c}-\tilde{p}_{c})^{2}}{2\omega(\tilde{p}_{c})^{2}}$
$=\frac{O\left(\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)^{1/2}\right)^{2}}{2\omega(\tilde{p}_{c})^{2}}$
$=O\left(\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\right)$
as $\mathrm{Var}\left(\sum_{c=1}^{C}Q_{c}\right)\to 0$, where $\omega_{c}(\tilde{p}_{c})\in[p_{c},\tilde{p}_{c}]\text{ or }[\tilde{p}_{c},p_{c}]$.
Figure C.1: Scatter plots of logit mean and variance pairs $(\mu_{c},\sigma_{c}^{2})$ on real models. We see that these seem to remain constrained to a compact set, with no much difference in the support as we increase the number of classes by an order of magnitude (from CIFAR-10 to CIFAR-100).
A key assumption in Theorem C.2 is the logit means and variances being restricted to a compact set. In Fig. C.1 we see empirically that this is approximately the case for HET and SNGP models, independently of the number of classes.
Theorem C.2 is of practical value as $M(\mathcal{K}):=\log\left(\frac{1+\delta/u}{1-\Delta}\right)$ is well-behaved: